# CAMeleon: Interactively Exploring Craft Workflows in CAD

### v1

@@ 8.1. Results
All participants succeeded in finding their desired workflow and built successfully, Figure 14 shows the resulting physical objects.
Figure 15, shows the results of the survey. Most participants found the software easy to use (10/12 gave it a score of more than 3/5), well-integrated (12/12 rated ¿3, and 8 gave it a 5/5), and not overly complex (11/12 rated this 1, 1/12 rated it 2). 10/12 participants indicated they would like to use this software frequently (2 gave it a neutral 3). Similarly, ease of use was rated positively by 10 participants and 2 gave it a neutral score.Most participants believed that people would quickly learn how to use the software (10 gave it a 5). The total mean SUS score amounted to 88.3.
On the more task-specific questions, 11/12 participants indicated the fabricated object met their expectations one was neutral on this. 8 participants indicated that they were confident in deciding their workflow, 2 gave this a negative score (2/5) and 2 were neutral. All participants indicated that they thought it was easy to change their workflow after seeing initial results and explored alternative processes.
@@ 8.2. Discussion
Positive User Experience Most participants found the software simple and easy to use, allowing those with no prior experience to get started smoothly. P1 highlighted that ”the workflow was nice, so I know what I’m doing, and how many steps are there,” appreciating the clear step-by-step guidance. P9 echoed this sentiment, stating, ”It is very direct. There is no complexity involved in it.” Similarly, P10 noted that ”the UI/UX design is pretty friendly to people like me,” even though they had little experience in fabrication. P6 also emphasized, ”I really love the detailed steps… because I don’t have previous (fabrication) experience. These steps helped me understand what to do, and made the whole process easy to follow.” These responses suggest that the software is well-designed for novice users, providing an accessible and straightforward navigation experience.
Workflow Comparison Users highlighted the visualization of workflows and comparison features as key strengths, helping them make informed decisions throughout the process. P2 noted, “I just want to see how it turns out, how it showed up, yeah, it’s more like aesthetic and a visualization,” emphasizing the importance of visual feedback in decision-making. Similarly, P5 shared, “So when I was doing that, it helped me to realize, okay, this is something that I really want to do. Maybe it was showing me a 3D model, which might be very close to how things would look like after. So it helped a lot.” P9 found the drag-and-drop comparison feature especially useful, stating, “I like how I can drag and drop the two or more workflows and, you know, compare them. I don’t think I am going to change any of it. It’s very self-explanatory and very distinguished software to use,” highlighting how this feature made it easier to select the desired production method. P8 echoed this sentiment, explaining that “the direct comparison of the workflow. It’s definitely helpful for me to find, like, finalize my choice.” Overall, participants appreciated the comparison feature for helping them understand the differences between various options, with visualization playing a critical role in decision-making.
Human-Centered CAM workflows Nearly all participants praised the step-by-step guidance, emphasizing how it helped them understand and follow the workflows. P10 highlighted, “I don’t have any previous experience, you know, having this. But the point is, you have the example for me, so which makes it super clear and explicit. It’s no matter the guidance on the software or the sample that you’re prepared for the users to, you know, anticipate what kind of results we really get.” Both P1 and P11 similarly appreciated this feature. P1 remarked, “This kind of handmade guidance. It’s usable. Yeah, it was super useful because I didn’t know what to do after.” Despite many users having limited or no prior experience in fabrication/craft, the detailed guidance provided by the software boosted their confidence in completing their projects.
Feature Requests: Expanding Support for Advanced Users Several participants noted that adding more advanced features could improve the experience for those with more expertise. One common suggestion was to include difficulty ratings and clearer pros and cons comparisons between workflows. For instance, P7 commented, “Maybe you could have a scale, like, ‘we advise you to use this one because it might be the easiest for what you want to achieve.” In the future, incorporating more advanced geometric analysis could further enhance the decision-making process for users. Similarly, P11 proposed using a star rating system or simple descriptions like “this workflow is suitable for beginners” or “this one saves time” to guide decision-making more effectively.
Several users highlighted the value of integrating generative AI functionality to simplify design and workflow selection. P6 suggested, “If you could embed that like a generative model in your software… that would be really helpful.” Incorporating AI-driven features or a chat-bot that recommends workflows based on user input could streamline the user experience, particularly for those who lack expertise in 3D modeling or fabrication.

### v2

@@ 9.2. Results
Figure 12 shows how each of the workflows implemented in CAMeleon allowed exported the resulting Stanford Bunny. Each of the workflows could be built up from modules existing in CAMeleon. Figure 13 outlines this reuse on a meta-level.
In re-implementing workflows, we used the Model Loading module to import STL files and prepare models for subsequent geometry processing. Each workflow involved distinct processing steps tailored to its goals:
For EscapeLoom, organic surfaces were converted into frame structures by adding warps parallel to surface curves and columns for closed surfaces. This was achieved using the Intersect module for dividing surfaces into frames and the Tube Gen module for creating volumetric structures. Fabrication was supported through the Export STL module.
In StackMold, 3D models were divided into planar contours to create layered molds. The Intersect module sliced the geometry and dealt with the bounding box into contours, while the Simplify Points/Lines refined inner paths and outer boundaries. The Loft module connected the contours into molds, and the Extrude to Direction module transformed them into volumes for Visualization, with cutting paths generated via the Export SVG module.
For Bamboo Agents, hollow frame structures were constructed by dividing models into triangular elements using the Hollow Triangle Gen module. Dimensions were adjusted with the Vertex Scaling module, and the Loft and Extrude to Direction modules were used to get the main structure and fixed structure. Fabrication was facilitated through the Export STL module.
Across all workflows, modules such as Model Information Visualization, 3D Visualization, and Tutorial & Info provided visual feedback, interactive guidance, and iterative support, ensuring a smooth and efficient design process. Figure 13 highlights the modules reused in adapting these workflows.
During the implementation process, we discovered that while the software behind these workflows handles geometric calculations, it falls short in accurately recording the full complexity of the workflow. Workflows still require tacit knowledge of craftsmanship. Key aspects of fabrication, which also rely on experiential knowledge, remain outside the scope of what these tools can document. While they help reduce certain challenges, they cannot fully represent the entire craft workflow, leaving gaps in the process that only skilled craftsmen can bridge. Such tacit knowledge is often overlooked or not documented in research papers. Once more, this calls for better documentation of Human-Centered CAM workflows. Capturing its full complexity would require adding tutorials and learning interfaces. Such implementation, while valuable, is beyond the scope of our current implementation and remain an area for future exploration.
10. User Evaluation: a Design Workshop
To understand the practical utility of CAMeleon for designers, and to observe its role in a design process, we conducted a design workshop in which teams of designers design and fabricate objects of their choice. Our objective was to study what role CAMeleon plays and how/whether it facilitates the intended interactive exploration of fabrication workflows.
Participants. We recruited 16 participants (11f/5m, 0 other/non-binary, mean age= 24.94 years, SD = 2.19) with varying levels of experience in fabrication—ranging from complete novices to advanced users (see Table 3).
ID | Age | Experience | Group | Background/Description |
P1 | 26 | novice | G1 | Law background; no fabrication experience. |
P2 | 27 | advanced | G2 | CS background; experienced in laser cutting and 3D printing. |
P3 | 23 | basic | G3 | CS background; took hardware courses (Human Robot Interaction). |
P4 | 24 | basic | G4 | CS background; took hardware courses (Human Robot Interaction). |
P5 | 28 | novice | G1 | Information Science background; no fabrication experience. |
P6 | 27 | novice | G2 | Technology and product design background; no fabrication experience. |
P7 | 23 | novice | G3 | CS background; no fabrication experience. |
P8 | 23 | advanced | G4 | Architect; skilled in CAD/CAM with 6 years of experience. |
P9 | 28 | advanced | G1 | Fabrication researcher; CS expertise. |
P10 | 26 | advanced | G2 | MBA background; experience in pottery, crochet, and wood cutting. |
P11 | 27 | basic | G3 | Digital artist; trained in CAD (C4D, Blender) and laser cutting. |
P12 | 26 | advanced | G4 | Digital artist; proficient in laser cutting, 3D printing, CNC, and ceramic printing. |
P13 | 24 | advanced | G1 | Architect; completed independent fabrication projects. |
P14 | 20 | basic | G2 | Maker space worker; basic knowledge of processes. |
P15 | 24 | advanced | G3 | 6 years of design experience. |
P16 | 23 | basic | G4 | Maker space worker; basic knowledge of processes. |
Tasks. Participants were assigned three tasks: (Task 1) brainstorm and discuss methods to fabricate models they were asked to prepare/select before the workshop; (Task 2) then they studied alternative fabrication workflows using provided objects (shown in Figure 14) fabricated using different workflows, and discussed challenges and considerations when switching workflows; and (Task 3) used CAMeleon to collaboratively design and fabricate objects.
Procedure. The workshop took place in two sessions: one plenary session with all participants, which took 105 minutes, and follow-up fabrication sessions with the individual teams of 4. The plenary workshop began with a pre-workshop demographics survey taking 10min. Participants were randomly assigned to 4 teams of 4, we ensured each group had a mixture of participants with different levels of expertise. Task 1 took place in small groups and took 20min. Between Task 1 and 2 we gave a brief presentation explaining what a full fabrication workflow looks like, at the example of 3D printing. Task 2 was conducted within the same teams took and 20min, followed by Task 3 which took 30min. The workshop concluded with about 15min of short presentations of the different teams in which they shared their final workflows, considerations and discussed the role of CAMeleon. We scheduled a follow-up session with each team for the actual fabrication. Participants were free to take breaks and communicate with other teams.
Data collection. The goal was to understand the role and usefulness of CAMeleon during the design process. Researchers took notes during the different tasks to capture participants’ decision-making processes, challenges, and comparison of workflows. Participants were given whiteboards and paper to make notes, which have been captured by researchers, as well as continuous video running during the tasks.
During the follow-up fabrication session, we engaged in reflective semi-structured qualitative interviews with participants to gather additional insights about their experiences, challenges, and perceptions of CAMeleon’s support in completing their workflows.
10.1. Results
All teams successfully designed and fabricated their intended artifacts in the workshop as shown in Figure 15. Based on the notes, observations, and interviews during the workshop we report the following findings:
10.1.1. Supporting Users with Varying Skill Levels and Backgrounds
We observed that level of expertise played a big role in how users engaged with the presented tasks and CAMeleon. Here we break that down along the same dimensions of expertise we highlighted in the demographics.
Novices and Entry-Level Users. Beginners often participated using rather simple ”volume models.” These straightforward models allowed novices to focus on following CAMeleon’s guidance step-by-step without overthinking each part of the fabrication process. Observations showed this simplicity encouraged gradual exploration and experimentation with different fabrication methods. We found that CAMeleon flattened the learning curve by offering comparisons of workflows and materials and providing realistic previews. P1 (novice) found the tool accessible and helpful: ”I think I would definitely use it if there is such a need, …I think it is quite clear… you put the file in, and then drag it to get a preview, and you can roughly know what the result is, and there are labels below, and you can roughly know what materials are, which is quite interesting.” P6 (novice) appreciated that the software provides detailed step-by-step processes: ”The software providing the model gives a step-by-step detailed process, and the ability to compare different models is quite helpful.” Initially, when choosing a workflow, P6 did not think about specifics like material preferences. Still, by interacting with the model, they were guided to consider features like depth, width, and structure. We found that participants learned about unknown fabrication workflows as they explored, as P6 mentioned CAMeleon:”When I worked with the model, it guided me to think about the specific features I was looking for. This helped me plan better…By comparing different models—observing details such as depth, width, distinct layers, and overall structure—it helped me decide which design was easier and more convenient to build.”
Before using the software, group discussions often showed ”mutual influence.” If an advanced user proposes to use 3D printing, the inexperienced members will usually be attracted by this ”expert idea” and will directly accept the solution. This situation is mainly because Novice and Basic users do not have enough knowledge of other processes and cannot independently evaluate, which leads them to rely on the opinions of advanced users. However, when CAMeleon was introduced, the behavior of participants changed. They all explored the software process individually, and then the users returned to the team with the fabrication methods they had selected or explored in the software for group discussion. The discussion at this stage was more efficient and diverse because each participant had certain background information and personal insights, and the discussion was no longer limited to the single idea of the expert but extended to a higher level of ideas around the pros and cons of different processes.
This also indicated that: as participants engaged with CAMeleon, their initial ambiguity about fabrication workflows dissipated. For instance, participants collectively explored the trade-offs between workflows during group discussions, debating options like laser cutting versus 3D printing for specific tasks. This collaborative process enabled them to articulate the steps involved, compare methods, and weigh the pros and cons of each. To be more specific, reflecting on what the participants initially imagined the workflows looked like, they lacked many details. For example, during task 1, Group 4 specified that they would 3D scan an object and ”draw horizontal and vertical lines” on the model to use the laser cutter and 2D shaper to create the final artifact. There is a considerable gap between this imagined workflow and what happens in the fabrication workflow. Detailing these nuances in an interface enables these participants to understand the trade-offs of each workflow better and mitigate some of the ”in-situ” challenges that come up with less thoughtful designs.
CAMeleon’s interactive guidance function helps users get familiar with the fabrication process and increase their confidence. However, the fabrication process still requires practical experience, and guidance alone cannot bypass all difficulties. For example, P1 is a law student who quickly learned the steps of human-centered CAM by reading the interactive guide when using CAMeleon. This made P1 feel that they had mastered the process. However, P1 repeatedly encountered difficulties when replicating the fabrication workflow. In this case, P1 returned to the software to view the interactive guide again and re-understood every detail in the steps. Through repeated comparison and practice, P1 gradually realized that the ignored details were not dispensable and directly affected the operation’s success.
Users with Basic Fabrication Experience. For participants with basic fabrication knowledge, CAMeleon bridged the gap between digital design and fabrication. P11 highlighted CAMeleon’s ability to expand creative possibilities: ”I definitely think it can help. The biggest inspiration for me lies in the materials and fabrication aspect. For example, 3D modeling can be done in tools like AutoCAD, C4D, or Blender, but when it comes to bringing something to life, without this software, I might just to using 3D printing with some default material to handle it. However, with this software, it has expanded my thinking and made me consider more possibilities.” P11 also mentioned ”[CAMeleon] let them make appropriate comparisons before fabrication.” P4, who has experience in robotic hardware design, described its potential in prototyping and reducing trial-and-error: “When we designed soccer robots, we made a rough prototype to confirm feasibility before refining materials and details. If software could help confirm the entire workflow at the start, it would be more efficient.”
A noteworthy example was P11’s use of CAMeleon to innovate process combinations. Initially, P11 used laser cutting to create an internal support structure and applied clay to shape it. Later, inspired by the soft texture of needle felting workflow by reintroducing the model to CAMeleon, and following the interactive guidance to learning steps, P11 added a needle-felted layer on top of the clay structure. By using CAMeleon to iterate and fine-tune these workflows, P11 completed a project with rich layers and unique textures. This highlights how CAMeleon supports not only traditional workflows but also innovative explorations that transcend material and process limitations.
Advanced Users. Advanced participants had experience with CAD tools like Rhino or Grasshopper. For them, CAMeleon acted as a complementary tool to explore fabrication possibilities. P8, an architecture professional with six years of experience, described CAMeleon as: ’For me, it’s more like a dictionary tool to see what’s possible, and then I can use other software to further refine it.’ P8 also proposed two hybrid usage modes for advanced users: 1) As an ”Interactive Dictionary” CAMeleon helps users explore workflows, compare options, and visualize results. This enables users to identify suitable workflows early and transition to their preferred design tools (e.g., Rhino, Fusion) to build the workflow and detailed adjustments. 2) For Workflow Comparison and File Preparation CAMeleon streamlines early steps, such as generating machine-ready files (e.g., SVG, STL), which users can then refine in CAD tools to meet project-specific design or fabrication needs. This avoids manual setup and repetitive adjustments. P12, a digital craft professional, highlighted its utility for installations and decision making: ”I think it can definitely support me. When it comes to comparing workflows, it’s actually quite useful. It might be even more beneficial for designers, as they often need to compare different materials. For other cases, like creating [art] installations, this kind of comparison would also be helpful.”
The exploratory potential of CAMeleon is further exemplified by P8’s group work during the workshop. P8, an architect, originally prepared a model of what was called a “candelabra” (see Figure 16a). At that stage, P8 did not have a clear idea of how to fabricate the model or a definitive final design scheme. By importing the model into CAMeleon—despite the tool primarily targeting volumetric models and the initial design not fully meeting those criteria—and browsing different fabrication workflows to understand the requirements and limitations, P8 began an iterative exploration process.
Based on the insights from CAMeleon regarding the model’s needs, P8 refined the design in successive steps, exploring various workflows. This process led to successfully processing the model in CAMeleon (see Figure 16b) and exporting a lasercut-ready SVG. P8 then imported the SVG into Rhino (see Figure 16c) for further refinement, iteratively carving out a central space to accommodate a candle (see Figure 16d). Through this process, CAMeleon enabled P8 and the team not only to engage in exploration in fabrication, but also to iteratively refine the design—using feedback to inform choices and leveraging CAMeleon to visualize and confirm those decisions. The final result is shown in Figure 15c.
10.1.2. Promoting Exploration in Fabrication
P11 reflected on how her initial fabrication approach was constrained by familiarity, leaning heavily toward 3D printing: ”…If I were to fabricate my own model, I might not know what materials to use. Initially, I tended to lean toward 3D printing because I was relatively more familiar with that manufacturing process. I didn’t know much about other fabrication methods… But after attending the workshop, I realized it’s possible to use other materials, like cotton thread, paper, or bamboo… to incorporate into the model. This was something I hadn’t considered before.” P4 described CAMeleon as a valuable reference of existing workflows: “It broadens your thinking and provides instructions… I learned new methods.” Similarly, P8 emphasized the tool’s role in listing fabrication workflows:“It lists multiple possible fabrication solutions, helping me explore workflows I’m not familiar with. At the very least, it gives me an idea of what those workflows might looks like.”
In the initial brainstorming period, participants were not just thinking about fabrication process but also about the final artifact and its materiality. For example, on one of the whiteboards, Group 2 had initially envisioned creating a dog or cat. Rather than relying on plastic/wood as the primary materials for these fabrication workflows that may be easier, they felt drawn to softer materials that dictated the workflows that they ended up using(needle felting) and CAMeleon suggested having foam and wool to create a less rigid body. The material selection played a significant role in the design process, they indicated a preference for workflows that matched their resulting vision of how the artifact would feel. CAMeleon visualizes this for them and enables them to explore which material fits their vision best.
10.1.3. Combining Computational Fabrication and Craft
Participants observed a gap between traditional craftsmanship and digital fabrication, which they think CAMeleon could bridge. P11 highlighted the divide:”…Everyone’s [fabrication] mindset is too digitalized… There’s either no one exploring, or there’s this clear divide between traditional [fabrication] and digital. I can feel it traditional methods are very traditional, with their own logic that’s deeply rooted in tradition, while digital methods are completely different…However, digital objects often become repetitive and lose the unique touch of traditional materials…There’s no exploration of something in between.” Reflecting on a hybrid model combining a copper coating , bamboo which made by CAMeleon, P11 remarked:: ”That process was very novel to me…It expands my mindset and lets me integrate traditional crafts with modern technology,” P1 noted “I think there is a bit of a gap [between craftsmanship and digital fabrication]” From P1’s perspective, one major benefit of human-centered-CAM workflows is that “it’s like the machine helps me create a framework, and I just need to add material on top.” This not only saves time and materials but also increases creative freedom: “There are certain structures I definitely wouldn’t be able to make if I were just doing by hand.”
During Group 2’s discussion, while brainstorming how to create a “fluffy dog,” the group collaboratively sketched a voluminous design, aiming to emphasize the artifact’s soft, furry texture. Although this artifact inherently aligns with traditional handcraft techniques, such as felting or crochet, the group, including advanced maker P10 (who has extensive experience in pottery, crochet, and wood cutting), expressed a preference for integrating ”digital fabrication tools”—specifically, using an electric clipper to shape the structure. This preference underscores how even advanced makers, despite their strong backgrounds in manual techniques, recognize the limitations of purely traditional approaches when faced with specific challenges like precision shaping or structural consistency.
10.1.4. Interface and Enhancements
Participants liked the simple and straightforward design of CAMeleon. P9 noted:The software quickly takes one object and shows it in different ways…It lets me compare different manufacturing workflows very clearly. Similarly, P3 highlighted the interface: ”…Each option has detailed instructions, which gives users a clear expectation of what they can achieve.” P1 explains that ”Overall, I think the software is quite straightforward” and emphasizes that “most of the interactive guidance is really helpful for a beginner like me.”
However, as this is research software, there were plenty of feature requests to improve CAMeleon:
Integration with CAD Tools P9 noted the lack of integration with design tools as a limitation: “It is a standalone page and not integrated into design tools, which makes it less convenient. It would be better if I could directly modify the model.” P12 suggested that integrating CAMeleon with modeling software could allow users to adjust models during the design process, rather than repeatedly revisiting designs after encountering manufacturing issues: ”If it could integrate with other design software, like modeling tools… it could allow users to adjust the model during the design process to some extent, rather than having to repeatedly modify the model after completing the design due to manufacturing-related issues.” P8 suggested integration with tools like Grasshopper and Rhino to enhance compatibility with established workflows: “It would be much more convenient if the webpage could directly connect with these commonly used software tools.”
Improving Performance In the workshop, some users were dissatisfied with the running speed of CAMeleon, which affected their experience to a certain extent. Modules in CAMeleon currently are there to support a large variety of design workflows, and are not yet optimized for performance. When loading or manipulating larger models, this sometimes caused latency problems.
Other suggestions P4 proposed incorporating large language models (LLMs) to enable natural language interaction, particularly for less experienced users: “Getting faster, more intuitive feedback through natural language could make software easier to use.”. P4 suggested a feature to recommend alternative modules when specific steps fail or are infeasible: “If a workflow has several modules and one of them doesn’t work, it would be helpful if the software could recommend new methods or replacements to adjust the workflow quickly.”. P1 pointed out that even if CAMeleon can generate machine instructions (e.g., SVG or CSV), some beginners might still struggle: “…may still not be able to use the machine. Like how to adjust the settings, or preparing the document and then putting it into an app… I don’t really understand those details.”. P11 suggested adding more examples of printing successes and failures, saying that it ”If you could include more details, like examples of printing failures and successes, it would make things clearer for users without much experience.” P12 highlighted the value of showing practical data like “budget or weight,” especially for larger objects,“One suggestion I made earlier was to include more practical data, like budget or weight. Sometimes, when objects get larger, different materials can make them quite heavy.”
During the workshop we find that, advanced participants often approached CAMeleon with a predetermined mindset regarding fabrication workflows for different parts during the model design stage. For instance, some users had clearly planned, “This part should be 3D printed, and that part should be laser cut.” Currently, CAMeleon treats the entire model as a single object, even when it contains multiple components, which means users cannot assign different fabrication workflows to individual parts.
Additionally, the software does not allow for differentiation within a single model when elements of varying scales exist. An example is P15, who brought a sand table architectural model consisting of floor slabs, beams, walls, tables, and chairs. Parameters optimized for fabricating the walls at 10 cm in height could not simultaneously achieve sufficient detail for the chairs at less than 3 cm. Without the ability to separate workflows for different parts of the model, P15 found it impossible to balance fidelity and efficiency. Faced with these limitations, P15 simplified their model by removing intricate details like tables and chairs, retaining only the primary structures such as floors and walls. This allowed them to experiment with alternative fabrication workflows that better suited the simplified model. Expert functionality, such as supporting assemblies or segmenting geometry, could address these challenges in the future.
@@ 10.2. Discussion
Overall, our evaluation showed that CAMeleon played a leading role in participant’s design processes. It served as a platform for exploration, comparison, and education around different fabrication workflows.
10.2.1. Bridging the Gap Between Guidance and Practical Application
We found that CAMeleon and the Human-Centered CAM workflows enables users without knowledge about specific workflows to compare different options and even fabricate using workflows they may initially have been unaware of. However, as P1’s case revealed: theoretical knowledge alone does not fully prepare users for the nuances and unexpected variables that arise during actual fabrication. This highlights the importance of tacit knowledge in fabrication workflows—knowledge that extends beyond replicating steps and often defies easy articulation, as ’we can know more than we can tell’ (Polanyi, 1967). Future iterations of CAMeleon could provide real-time, contextual support, such as AR or voice-based guidance, to bridge this gap and help users navigate the dynamic and emergent nature of craft workflows.
10.2.2. Bridging Traditional Craftsmanship and Digital Fabrication
CAMeleon enables users to create works that retain the unique tactile qualities of handcraft while introducing innovative design possibilities. The introduction of Human-Centered CAM crucially extended this beyond what a machine can do, but explicitly embraces human effort in the export format. By offering these at the same level of fidelity, the workshop shows CAMeleon allowed designers to focus on personalized exploration. Even advanced users with expertise in fabrication embraced virtual comparisons, illustrating that hybrid workflows benefit both beginners and experienced makers. Moreover, it showed CAMeleon empowers rather than replaces traditional skills, fostering creativity and innovation such as remixing existing workflows (Zoran and Buechley, 2013).
10.2.3. Enhancing Integration in the Design-to-Fabrication Workflow
Advanced users of CAMeleon revisit the platform to review guidance before returning to their CAD tools to adjust designs. This pattern suggests a need for stronger support in the feedback loop between design and fabrication. Participants pointed this out as a clear opportunity for progress, albeit outside of the scope of this paper.
We see such tighter integration between CAMeleon and design environments as a fruitful step forwards, such integration could expose constraints of fabrication workflows in the design process and maybe nudge designers towards alternative fabrication workflows as they make decisions on their design. Generally, crossing the divide between CAD and CAM more naturally may be a powerful direction for the development of future tools.
10.2.4. Rethinking Workflow Choices Through Collective Exploration
Our observations show that workflow decisions are often influenced by experienced participants. By democratizing access to workflows, CAMeleon changed this pattern. This gave inexperienced users confidence to return to group discussions with their own insights, making the conversations more diverse and productive. The impact of CAMeleon goes beyond just individual exploration. It changed how participants viewed their roles in decision-making, allowing them to confidently challenge the status quo. This shifted the focus of discussions in the design process to consider a wider range of perspectives.
These findings suggest a potential for improving interdisciplinary collaboration. In many hardware design projects, each member is often working within the boundaries of their own expertise. Tools like CAMeleon could encourage experts from different fields to step outside of their specific expertise, fostering shared insights and more integrated problem-solving. This approach has the potential to create a collaborative environment where decisions are enriched by contributions from all perspectives, leading to more innovative and well-rounded outcomes.